# Fractional Semilinear Equations on Hyperbolic Spaces

### v1

@@ 4.2. Critical case $p=\frac{n+2s}{n-2s}$
$u(x)=\int_{B_{R}(o)}G(\rho(x,y))u(y)^{p}\,dV_{y}+\int_{\mathbb{H}^{n}\setminus B_{R}(o)}G(\rho(x,y))u(y)^{p}\,dV_{y}.$
For the local part, we have that for any fixed $y\in B_{R}$, $\rho(x,y)\to\infty$ as $\rho(x,o)\to\infty$. By kernel estimate, $G(\rho(x,y))\lesssim\rho^{-1+s}e^{-(n-2s)\rho(x,o)}$. Hence
$\int_{B_{R}}G(\rho(x,y))u(y)^{p}\,dV_{y}\lesssim\rho^{-1+s}e^{-(n-2s)\rho(x,o)}\int_{B_{R}}u(y)^{p}\,dV_{y}\to 0$
@@
$\left|\int_{\mathbb{H}^{n}\setminus B_{R}(o)}G(\rho(x,y))u(y)^{p}\,dV_{y}\right|\leqslant\|G(\rho(x,\cdot))\|_{L^{a^{\prime}}(\mathbb{H}^{n}\setminus B_{R})}\;\|u^{p}\|_{L^{a}(\mathbb{H}^{n}\setminus B_{R})}.$
@@
$\int_{\mathbb{H}^{n}\setminus B_{R}}G(\rho(x,y))^{a^{\prime}}\,dV_{y}\lesssim\int_{R}^{\infty}\rho^{-1+s}e^{-a^{\prime}(n-2s)\rho}e^{(n-1)\rho}\,d\rho=\int_{R}^{\infty}\rho^{-1+s}e^{-(a^{\prime}(n-2s)-(n-1))\rho}\,d\rho$
We are able to choose $a$ sufficiently close to $1$, so that $a^{\prime}(n-2s)-(n-1)>0$. Thus, $\|G(\rho(x,\cdot))\|_{L^{a^{\prime}}(\mathbb{H}^{n}\setminus B_{R})}$ can be made arbitrarily small for large enough $R$. The proof is now complete. ∎
@@
5. Various Maximum Principles on $\mathbb{H}^{n}$

### v2

@@ 4.2. Critical case $p=\frac{n+2s}{n-2s}$
$u(x)=\int_{B_{R}(o)}G_{s}(\rho(x,y))u(y)^{p}\,dV_{y}+\int_{\mathbb{H}^{n}\setminus B_{R}(o)}G_{s}(\rho(x,y))u(y)^{p}\,dV_{y}.$
For the local part, we have that for any fixed $y\in B_{R}$, $\rho(x,y)\to\infty$ as $\rho(x,o)\to\infty$. By kernel estimate, $G_{s}(\rho(x,y))\lesssim\rho^{-1+s}e^{-(n-2s)\rho(x,o)}$. Hence
$\int_{B_{R}}G_{s}(\rho(x,y))u(y)^{p}\,dV_{y}\lesssim\rho^{-1+s}e^{-(n-2s)\rho(x,o)}\int_{B_{R}}u(y)^{p}\,dV_{y}\to 0$
@@
$\left|\int_{\mathbb{H}^{n}\setminus B_{R}(o)}G_{s}(\rho(x,y))u(y)^{p}\,dV_{y}\right|\leqslant\|G_{s}(\rho(x,\cdot))\|_{L^{a^{\prime}}(\mathbb{H}^{n}\setminus B_{R})}\;\|u^{p}\|_{L^{a}(\mathbb{H}^{n}\setminus B_{R})}.$
@@
$\int_{\mathbb{H}^{n}\setminus B_{R}}G_{s}(\rho(x,y))^{a^{\prime}}\,dV_{y}\lesssim\int_{R}^{\infty}\rho^{-1+s}e^{-a^{\prime}(n-2s)\rho}e^{(n-1)\rho}\,d\rho=\int_{R}^{\infty}\rho^{-1+s}e^{-(a^{\prime}(n-2s)-(n-1))\rho}\,d\rho$
We are able to choose $a$ sufficiently close to $1$, so that $a^{\prime}(n-2s)-(n-1)>0$. Thus, $\|G_{s}(\rho(x,\cdot))\|_{L^{a^{\prime}}(\mathbb{H}^{n}\setminus B_{R})}$ can be made arbitrarily small for large enough $R$. The proof is now complete. ∎
@@
5. Several Maximum Principles on $\mathbb{H}^{n}$
@@
Acknowlegment
The author would like to thank Professor Bruno, T and Professor Papageorgiou, E for their valuable comments and for informing some crucial recent results in this topic.